a savings account investment of $1400 at 9% coumpounded yearly for 6 months what is the future value?
Answer by stanbon(75887)

a savings account investment of $1400 at 9% coumpounded yearly for 6 months what is the future value?
----
A(t) = P(1 + (r/n))^(nt)
---
A(1/2) = 1400(1 + (0.09/1))^(1/2)
---
A(1/2) = 1500(1.09)^(1/2) = 1500*1.044 = $1566.05
===========
Cheers,
Stan H.
===========